OpenStudy (anonymous):

What is the equation of a line passing through (-8, -3) and having a slope of -4?

OpenStudy (anonymous):

choices: y = −4x + 20 y = 4x + 20 y = −4x – 35 y = 4x – 35 y = -4x + 4

OpenStudy (anonymous):

y = −4x + 20 y = 4x + 20 y = −4x – 35 y = 4x – 35 y = -4x + 4

OpenStudy (anonymous):

The equation of line having \((x_1, y_1)\) point and given slope is : \[y - y_1 = m(x-x_1)\] Here, \((x_1,y_1) = (-8,-3)\) and \(m = -4\), just plug in the values..

OpenStudy (anonymous):

\[y - (-3) = -4(x - (-8)) \implies y + 3 = -4(x+8)\]

OpenStudy (anonymous):

Solve right hand side by using Distributive property ie, distribute \(-4\) to the brackets and tell me what did you get?

OpenStudy (anonymous):

y = −4x + 20

OpenStudy (anonymous):

No, how you got that?

OpenStudy (anonymous):

Can you show me your steps?